OpenStudy (anonymous):

solve...-21le3x-4le-3

OpenStudy (anonymous):

sorry i didnt know how to make the less than equal to sign in the room

OpenStudy (anonymous):

\[21\le3x-4\le-3\]

OpenStudy (anonymous):

21≤3x−4≤−3 \[25\le3x\le1\] \[25/3\ge x\ge 1/3\]

OpenStudy (anonymous):

Dividing flips the signs.

OpenStudy (a_clan):

dividing with negative flips

OpenStudy (anonymous):

It doesn't work the other way.

OpenStudy (anonymous):

ok im confused lol

OpenStudy (anonymous):

i do think that the signs are flipped correct?

OpenStudy (anonymous):

A_clan is correct my last step should have the opposite signs. The result is not true, unless I'm missing something.

OpenStudy (anonymous):

@A_clan, thanks for pointing that out.

OpenStudy (a_clan):

The correct eqn 25/3 <= x <= 1/3 solution (-infinity, 1/3] U [25/3, infinity)

OpenStudy (anonymous):

so if i put this in set notation what would it be \[\left\{ x \left|25 \right| \right\}\]

OpenStudy (anonymous):

ok i messed that up

OpenStudy (anonymous):

@ A_clan That's an empty set though.

OpenStudy (a_clan):

its union

myininaya (myininaya):

so does everyone understand that if we have 3>2 then this inequality still holds if we divide both sides by 2 3/2>2/2

OpenStudy (anonymous):

wait for the set they give me to fill in it looks like this \[\left\{ x \left| ?lex \le? \right| \right\}\]

OpenStudy (anonymous):

I think ive lost myself lol

OpenStudy (anonymous):

the signs are going like this\[\le\]

OpenStudy (anonymous):

@ A_clan: your right sorry. @ myininaya, Yes(at least i do). @tena32, sorry for my mistakes and the confusion they have caused.

myininaya (myininaya):

if we divide by -2 then inequality flips we have 3>2 3/(-2)<2/(-2)

OpenStudy (anonymous):

thats ok, I make mistakes as well

OpenStudy (anonymous):

@ myinanaya: we were dividing by positive 3

OpenStudy (anonymous):

im trying to show yall how the solution set is set up but the equation thing isnt working for me

OpenStudy (a_clan):

solution in set-builder form: {x | x belongs to REAL Number, x>= 25/3 and x<= 1/3}

myininaya (myininaya):

i was trying to give an example of when to flip the inequality sign

OpenStudy (anonymous):

@myininaya thanks

OpenStudy (a_clan):

use appropriate symbol for "belongs to" and "Real Number" in set-builder form

OpenStudy (anonymous):

\[\left\{ x \left| -2\le x \le-2 \right| \right\}\]

OpenStudy (anonymous):

ok finally. is this how is is?

OpenStudy (anonymous):

{\[x|x \in \mathbb{R} , x \ge 25/3 \cup x \le 1/3\]}

OpenStudy (a_clan):

kd8cpk gave the correct notation for your answer

OpenStudy (anonymous):

ok i dont know if i have gotten yall confused but in the answer key they gave me the symbols are both \[\le\] they look like they didnt change the symbol. or am i just confused

OpenStudy (anonymous):

look at my 3rd post up. that is they way they have it set up for me to fill in my answer.

OpenStudy (a_clan):

25/3 <= x <= 1/3 you can also use this notation in your solution for conformity with the answer key. We gave you the other one, because it was easier for understanding

OpenStudy (anonymous):

ok so 25/3 and 1/3 are the fractions

OpenStudy (a_clan):

yes

OpenStudy (anonymous):

i got ya now. thanks!